# Spectral Criteria for the Asymptotics of Local Functionals of Gaussian Fields and Their Application to Nodal Volumes

### v2

@@ 1.4 Asymptotics of the fourth chaos
We showed in the previous paragraph that regular random waves, and in particular isotropic random waves, are subject to the general cancellation phenomenon of Theorem 1.4. In this section, we derive the exact variance asymptotics of the nodal volume of the isotropic random wave model, following the heuristic outlined after Theorem 1.8. The result we obtain encompasses and generalizes many findings in the literature concerning such asymptotics, which have primarily focused on dimensions $2$ and $3$.
The case $d=2$ is well-studied, with significant results found in works such as [Cam19, Kri13, Nou19] across various settings (spherical harmonics, arithmetic random waves, planar random waves). The exact asymptotics of the variance of the nodal length have been explicitly computed : there exists an explicit positive constant $C$ such that
@@
and $Z_{\lambda}$ satisfies a CLT. When $d=3$, it has been proved in [Dal19] the existence of a positive constant $C$ such that
@@
as well as a CLT. The next theorem proves a general version of these two facts. {mdframed}[backgroundcolor=white]
@@
Let $Y$ be the isotropic random wave model defined in introduction. Then there is a positive constant $C_{d}$ such that as $\lambda$ goes to +$\infty$+ ∞
@@ 3.2 Positive Hermitian measure and Bochner Theorem
Let $\Omega:V\rightarrow\mathcal{S}q(W)$. The two propositions are equivalents.
@@
If $W$ is a real vector space instead of a complex vector space, we can embed it in its complexification $W_{\mathbb{C}}=W\otimes\mathbb{C}$ via the mapping $w\mapsto w\otimes 1$. A map $\Omega:V\rightarrow(W^{*})^{\otimes 2}$ can be seen as a map taking values in $\operatorname{Sesq}(W_{\mathbb{C}})$ via this embedding, and one can extend the notion of positive definiteness for such bilinear maps $\Omega$. The conclusion of Lemma 3.4 then holds if one replace $(2)$ by saying that $\Omega$ is the Fourier transform of a (symmetric) positive Hermitian measure over $W_{\mathbb{C}}$. In particular, if $X:V\rightarrow W$ is a stationary Gaussian process with continuous covariance function $\Omega$, then its spectral measure $\mu=\widehat{\Omega}$ is a symmetric positive Hermitian measure over $W_{\mathbb{C}}$.
@@ 4.1 The nodal measure of random fields
4.2 Chaotic expansion of the nodal measure
@@
and one can apply the general formalism developed in Section 2.
Remark 4.2.
The zero set of $Y$ is independent of the choice of scalar product on $U$, and one can assume that $Y(0)$ is a standard Gaussian random variable. Similarly, a linear change of variables on $V$ modifies the nodal measure $dZ$ by a factor related to the determinant of the Jacobian transformation. One can assume that, for any fixed $u\in U$, $\nabla\langle Y(0),u\rangle$ is a standard Gaussian vector. Thus, the choice of scalar product on $V$ is arbitrary. However, one cannot assume that $X(0)=(Y(0),\nabla Y(0))$ is a standard Gaussian random variable. This is because the scalar product induced by $X(0)$ on $W=U\otimes(\mathbb{R}\times V)$ is generally different from the one inherited from any choice of scalar products on $U$ and $V$ due to a dimension argument. We must work with a rescaled version of $X$ and $f$ such that $X(0)$ becomes a standard Gaussian vector.
The semi-definiteness of $r$ implies that $r(-x)=r(x)^{T}$. Differentiating with respect to a vector $v$ at the point $x=0$, we find that $\operatorname{Cov}(Y(0),\nabla Y(0))$ is an anti-symmetric bilinear map. If $k=1$, this implies that $Y(0)\perp\nabla Y(0)$, and one can further assume that $X(0)$ is a standard Gaussian random variable, as explained in the previous paragraph. This fact partially explains why providing sufficient conditions to ensure the positivity of the variance is a more delicate task in the case $k>1$ and is difficult to treat in full generality. We show that, under natural additional constraints on the covariance function of the field, one can obtain asymptotic results. To compute the chaotic expansion explicitly, we use the following general lemma. {mdframed}[backgroundcolor=white]
@@
In light of Theorem 1.4 for the second chaos, we compute the spectral measure $\mu$ of the process $X=(Y,\nabla Y)$, and some quantities related to the second chaotic component $f_{2}$ of the tempered distribution $f$ defined in (4.1). {mdframed}[backgroundcolor=white]
@@
The first point follows from the fact that the Fourier transform derivation into multiplication. In particular, for $q\geq 0$ the Fourier transform of the differential $D^{q}r$, taking values in $U\times V^{\otimes q}$, is given when defined by $\psi\otimes(2i\pi\xi)^{\otimes q}$. As for the second point, notice that the tensor $(1,2i\pi\xi)\otimes(1,-2i\pi\xi)$ is of rank one and its image is exactly the line directed by the vector $(1,2i\pi\xi)\in\mathbb{R}\times V$. The conclusion follows directly from the definition of the image of $\mu$, since $\mu\ll\psi$. ∎
We now compute quantities related to the second chaotic component $f_{2}$ of the formal function $f:(x,D)\mapsto\delta_{0}(x)J(D)$. Let $\rho(0)$ be the density at 00 of the vector $Y(0)$. One can assume, as explained in Remark 4.2, that $Y(0)$ is a standard Gaussian vector, and that for one fixed vector $v\in U$, the vector $Y_{v}=\nabla\langle Y(0),v\rangle$ is a standard Gaussian vector. We define the following quantities for a vector $w$ perpendicular to $v$, that are associated to the chaotic projection of order $0,1,2$ and $4$ of the volume of a $u$-level.
$q=0$ :
$\alpha=\mathbb{E}[f(X(0))]$
$q=1$ :
$\delta=\mathbb{E}[f(X(0))Y(0)],\quad\chi=\mathbb{E}[f(X(0))\nabla Y(0)]$
$q=2$ :
$\Lambda=\mathbb{E}[f(X(0))[Y(0)Y(0)^{T}-\mathrm{Id}]],\quad\Theta=\mathbb{E}[f(X(0))Y(0)\nabla Y(0)]$
$\beta_{u}=\frac{1}{d}\mathbb{E}[f(X(0))(\|Y_{v}\|^{2}-d)],\quad\kappa_{vw}=\mathbb{E}[f(X(0))Y_{v}^{T}Y_{w}],$
$q=4$ :
$\zeta_{v}=\frac{1}{d}\mathbb{E}[f(X(0))(\|Y_{v}\|^{2}-d)[Y(0)Y(0)^{T}-\mathrm{Id}]].$
@@
$\alpha\neq 0.$
$\delta=\alpha u,\quad\chi=0.$
$\quad\Theta=0,\quad\text{and}\quad\Lambda=\alpha(uu^{t}-\mathrm{Id}).$
If $Y_{v}$ is independent to $Y_{w}$, for $w\perp v$ then
$\kappa_{vw}=0,\quad\beta_{v}=\frac{\alpha}{d}\quad\text{and}\quad\zeta_{v}=-\frac{\alpha}{d}\mathrm{Id}.$
@@
The firsts statements are direct consequences of the symmetry of the Jacobian map, and Lemma 4.3. The second part of the proof borrows a nice trick from [Not21]. Notice first that
$\alpha=\rho(u)\mathbb{E}[J(\nabla Y(0))].$
Then the law of the vector $Y_{v}$ is invariant by rotation, which implies that
$\|Y_{v}\|\perp\frac{Y_{v}}{\|Y_{v}\|}.$
Moreover, the formula for the moments of a $\chi^{2}$ distribution implies that
$\frac{\mathbb{E}[\|Y_{v}\|^{3}]}{\mathbb{E}[\|Y_{v}\|]}=d+1.$
Now
$\beta_{v}$ $=\frac{1}{d}\rho(u)\mathbb{E}[J(Y_{v},(Y_{w})_{w})\|Y_{v}\|^{2}]-\alpha$
@@
$=\frac{1}{d}\rho(u)\mathbb{E}\left[J\left(\frac{Y_{v}}{\|Y_{v}\|},(Y_{w})_{w}\right)\|Y_{w}\|^{3}\right]-\alpha$
$\overset{\perp}{=}\frac{1}{d}\rho(u)\mathbb{E}[J(Y_{v},(Y_{w})_{w})]\frac{\mathbb{E}[\|Y_{v}\|^{3}]}{\mathbb{E}[\|Y_{v}\|]}-\alpha$
$=\frac{d+1}{d}\alpha-\alpha$
$=\frac{\alpha}{d}.$
At last, by independence of $Y$ and $\nabla Y$,
$\zeta_{v}=\frac{\beta_{v}\Lambda}{\alpha}=\frac{\alpha}{d}(uu^{t}-\mathrm{Id}).$

### v3

@@ 1.4 Asymptotics of the fourth chaos
We showed in the previous paragraph that regular random waves are subject to the general cancellation phenomenon of Theorem 1.4. In this section, we derive the exact variance asymptotics and a CLT for the nodal volume of regular symmetric random waves, following the heuristic outlined after Theorem 1.4. The result we obtain encompasses and generalizes many findings in the literature concerning such asymptotics, which have primarily focused on dimensions $d=2,3$ for the isotropic random wave model. We also prove with the same method the exact variance asymptotics and the CLT for the number of critical points of real-valued isotropic random waves, which is a new result in the literature.
The case $d=2$ for the nodal volume of isotropic random waves is well-studied, with significant results found in works such as [Cam19, Kri13, Nou19] across various settings (spherical harmonics, arithmetic random waves, planar random waves). The exact asymptotics of the variance of the nodal length have been explicitly computed : there exists an explicit positive constant $C$ such that
@@
and $Z_{\lambda}$ satisfies a CLT. When $d=3$, it has been proved in [Dal21] the existence of a positive constant $C$ such that
@@
as well as a CLT. The next theorem proves a general version of these two facts, and includes the case of critical points of a real-valued isotropic random wave. {mdframed}[backgroundcolor=white]
@@
Assume either $Y$ is a symmetric regular random wave, or that $Y$ is the gradient of a real-valued isotropic random wave. Then there is a positive constant $C_{d}$ such that as $\lambda$ goes to +$\infty$+ ∞
@@ 3.2 Positive Hermitian measure and Bochner Theorem
Let $\Omega:V\rightarrow\operatorname{Sesq}(W)$. The two propositions are equivalents.
@@
If $W$ is a real vector space instead of a complex vector space, we can embed it in its complexification $W_{\mathbb{C}}=W\otimes\mathbb{C}$ via the mapping $w\mapsto w\otimes 1$. A map $\Omega:V\rightarrow(W^{*})^{\otimes 2}$ can be seen as a map taking values in $\operatorname{Sesq}(W_{\mathbb{C}})$ via this embedding, and one can extend the notion of positive definiteness for such bilinear maps $\Omega$. The conclusion of Lemma 3.4 then holds if one replace $(2)$ by saying that $\Omega$ is the Fourier transform of a positive Hermitian measure over $W_{\mathbb{C}}$. In particular, if $X:V\rightarrow W$ is a stationary Gaussian process with continuous covariance function $\Omega$, then its spectral measure $\mu=\widehat{\Omega}$ is a symmetric positive Hermitian measure over $W_{\mathbb{C}}$.
@@ 4.1 The nodal measure of random fields
4.2 Chaotic expansion
@@
and one can apply the general formalism developed in Section 2. To compute the chaotic expansion explicitly, we use the following general lemma. {mdframed}[backgroundcolor=white]
@@
We state one more lemma that we will use later in the proof. {mdframed}[backgroundcolor=white]
Lemma 4.3.
Let $X$ a standard Gaussian vector of size $n$ and $G$ be a function on $\mathbb{R}^{n}$ such that for $c\geq 0$ and $m$ an integer such that $G(cX)=c^{m}G(X)$. Then
$\mathbb{E}\left[G(X)\left(\|X\|^{2}-\mathbb{E}[\|X\|^{2}]\right)\right]=m\mathbb{E}[G(X)].$
Proof.
By the formula for the moments of a $\chi^{2}$ distribution, one has
$\frac{\mathbb{E}\left[(\|X\|^{2}-\mathbb{E}[\|X\|^{2}])\|X\|^{m}\right]}{\mathbb{E}[\|X\|^{m}]}=m.$
Using the identity $G(X/\|X\|)=G(X)/\|X\|^{m}$ and the independence between $X/\|X\|$ and $\|X\|$, we deduce that
$\mathbb{E}\left[G(X)\left(\|X\|^{2}-\mathbb{E}[\|X\|^{2}]\right)\right]$ $=\mathbb{E}[G(X)]\frac{\mathbb{E}\left[(\|X\|^{2}-\mathbb{E}[\|X\|^{2}])\|X\|^{m}\right]}{\mathbb{E}[\|X\|^{m}]}$
$=m\mathbb{E}[G(X)]$
We now compute quantities related to the second chaotic component $f_{2}^{(u)}$ of the formal function $f^{u}:(x,D)\mapsto\delta_{u}(x)J(D)$. We explicit some of the quantities associated to the chaotic projection of order $0,1,2$ and $4$ of the nodal volume of $Y$. Let $\rho$ be the density of the standard Gaussian vector $Y(0)$. We define the positive quantity
$\alpha_{u}=\rho(u)\mathbb{E}[J(\nabla Y(0))],$
and we will omit the index $u$ if $u=0$.
4.2.1 Chaotic expansion under hypothesis (H1)
In light of Theorem 1.4 for the limit variance of the second chaos, we compute the spectral measure $\mu$ of the process $X=(Y,\nabla Y)$, and some quantities related to the second chaotic component $f_{2}^{u}$ of the tempered distribution $f^{u}$ defined in (4.1) under hypothesis (H1). We recall that the process $X$ takes values in the space
$W=U\otimes(\mathbb{R}\times V).$
{mdframed}
[backgroundcolor=white]
@@ 4.2.1 Chaotic expansion under hypothesis (H1)
The first point follows from the fact that the Fourier transform derivation into multiplication. In particular, for $q\geq 0$ the Fourier transform of the differential $D^{q}r$, is given when defined by $\psi\otimes(2i\pi\xi)^{\otimes q}$. As for the second point, notice that the tensor $(1,2i\pi\xi)\otimes(1,-2i\pi\xi)$ is of rank one and its image is exactly the line directed by the vector $(1,2i\pi\xi)$. The conclusion follows directly from the definition of the image of $\mu$, since $\mu\ll\psi$. ∎
@@
Under hypothesis (H1) the second chaotic component $f_{2}^{u}$ has the expression, in an orthogonal basis on $W$
$f_{2}^{u}=\alpha_{u}\left[\mathrm{Id}_{U}\otimes\begin{pmatrix}[c|c]-1&0\\ \hline\cr 0&\frac{1}{d}\,\mathrm{Id}_{V}\end{pmatrix}+(uu^{T})\otimes\begin{pmatrix}[c|c]1&0\\ \hline\cr 0&0\end{pmatrix}\right].$
@@
$Y(0)$ and $\nabla Y(0)$ are independent. Combined with the symmetry of the Jacobian map and Lemma 4.2, we deduce that
$\mathbb{E}[f^{u}(X(0))Y(0)\nabla Y(0)]=0\quad\text{and}\quad\mathbb{E}\left[f^{u}(X(0))(Y(0)Y(0)^{T}-\mathrm{Id}_{U})\right].$
The invariance by rotation of the Jacobian map implies the existence of a constant $c$ such that
$\mathbb{E}\left[f^{u}(X(0))\left(\nabla Y(0)\otimes\nabla Y(0)^{T}-\mathrm{Id}_{U}\otimes\mathrm{Id}_{V}\right)\right]=c\mathrm{Id}_{U}\otimes\mathrm{Id}_{V}.$
The constant $c$ can be explicitly computed thanks to Lemma 4.3, since
$c=\frac{1}{dk}\mathbb{E}\left[f^{u}(X(0))\left(\|\nabla Y(0)\|_{2}^{2}-\mathbb{E}[\|\nabla Y(0)\|_{2}^{2}]\right)\right]=\frac{\mathbb{E}[f^{u}(X(0))]}{d}=\frac{\alpha_{u}}{d}.$
The conclusion follows. ∎
The fourth chaotic component $f_{4}$ of $f$ can be seen as a symmetric bilinear map on $S^{2}(W^{*})$, which can be identified as the direct sum
$S^{2}(W^{*})=S^{2}(U)\oplus S^{2}(U\otimes V)\oplus\left(U\otimes(U\otimes V)\right).$
{mdframed}
[backgroundcolor=white]
Lemma 4.6.
The subspace $U\otimes(U\otimes V)$ of $S^{2}(W^{*})$ is stable by $f_{4}$. Moreover, the restriction of $f_{4}$ to this subspace, denoted by $f_{4}^{r}$, has expression
$f_{4}^{r}=-\frac{\alpha}{d}\mathrm{Id}_{U}\otimes\left(\mathrm{Id}_{U}\otimes\mathrm{Id}_{V}\right).$
Proof.
The fact that the subspace $U\otimes(U\otimes V)$ directly follows from the symmetry of the Jacobian map. The restriction of $f_{4}^{r}$ to this subspace is equal to
$f_{4}^{r}$ $=\mathbb{E}\left[f(X(0))\left(Y(0)Y(0)^{T}-\mathrm{Id}_{U}\right)\otimes\left(\nabla Y(0)\otimes\nabla Y(0)^{T}-\mathrm{Id}_{U}\otimes\mathrm{Id}_{V}\right)\right]$
$=-\mathrm{Id}_{U}\otimes\mathbb{E}\left[f(X(0))\left(\nabla Y(0)\otimes\nabla Y(0)^{T}-\mathrm{Id}_{U}\otimes\mathrm{Id}_{V}\right)\right]$
$=-\frac{\alpha}{d}\mathrm{Id}_{U}\otimes\left(\mathrm{Id}_{U}\otimes\mathrm{Id}_{V}\right),$
where the last line follows from Lemma 4.5. ∎
4.2.2 Chaotic expansion under hypothesis (H2)
Under hypothesis (H2), in a good choice of scalar product on $U$ and $V$ one can write $Y=(Y_{1},Y_{2})$, where $Y_{1}\perp Y_{2}$, and $Y_{1}$ is a real valued Gaussian field such that $\operatorname{Var}(Y(0),\nabla Y(0))=\mathrm{Id}$. Since $(Y_{2},\nabla Y_{2}(0))$ is not assumed to be a standard Gaussian vector, one must work with a linear transformation of the process $(Y_{2},\nabla Y_{2})$, so that it is point-wise a standard Gaussian vector. {mdframed}[backgroundcolor=white]
Lemma 4.7.
Under hypothesis (H2), the spectral measure is block diagonal, where the first block corresponds to the spectral measure of $(Y_{1},\nabla Y_{1})$, and the second block to a linear transformation of the spectral measure of $(Y_{2},\nabla Y_{2})$. In particular, the first block has the expression
$\mu_{1}=\psi_{1}\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
Proof.
The block diagonal property is a direct consequence of the independence, and the expression of the spectral measure of the first bloc follows directly from Lemma 4.4. ∎
{mdframed}
[backgroundcolor=white]
Lemma 4.8.
Under hypothesis (H2), the second chaotic projection $f_{2}$ is block diagonal, and the first block has expression
$f_{2,1}=\alpha\begin{pmatrix}[c|c]-1&0\\ \hline\cr 0&M\end{pmatrix},$
where $M$ is a matrix with unit trace.
Proof.
The block diagonal property is a direct consequence of the independence of $Y_{1}$ and $Y_{2}$ and the symmetry of the Jacobian map. Moreover, one has
$\mathbb{E}[f(X(0))Y_{1}(0)\nabla Y_{1}(0)]=0\quad\text{and}\quad\mathbb{E}\left[f(X(0))(Y_{1}(0)^{2}-1)\right]=-\alpha.$
Let
$M=\frac{1}{\alpha}\mathbb{E}[f(X(0))(\nabla Y_{1}(0)\nabla Y_{1}(0)^{T})-\mathrm{Id}_{V})].$
Then
$\operatorname{Tr}(M)=\frac{1}{\alpha}\mathbb{E}\left[f(X(0))\left(\|Y_{1}(0)\|^{2}-\mathbb{E}[\|Y_{1}(0)\|^{2}]\right)\right]=1,$
where the last equality follows from Lemma 4.3 applied to the determinant as a function of the first column. ∎
4.2.3 Chaotic expansion under hypothesis (H3)
Under hypothesis (H3), one has $(Y,\nabla Y)=(\nabla F,\nabla^{2}F)$, where $\nabla^{2}F$ denotes the Hessian of the random field $F$, which takes values in $S^{2}(V^{*})$, the space of symmetric bilinear map on $V$ equipped with the Frobenius scalar product. We denote by $q$ the covariance function of $F$ and $\omega$ its spectral measure. In that case, the process $(\nabla F,\nabla^{2}F)$ takes values in the space
$W=V\times S^{2}(V^{*}).$
On can assume that the scalar product defined on $V$ is such that $\nabla F$ is point-wise a standard Gaussian vector, and after rescaling, that $F$ is point-wise a standard Gaussian vector. In general $\nabla^{2}F$ is not a standard Gaussian vector, and one must work with a linear transformation, denoted by $TF$, such that $TF$ is a standard Gaussian vector on $\operatorname{Sym}(V)$. Note that the distribution of $TF$ is exactly $\frac{1}{\sqrt{2}}\mathrm{GOE}$, where $\mathrm{GOE}$ is the Gaussian Orthogonal Ensemble. If $F$ has an isotropic distribution (i.e. $q$ is a radial function), then the linear transformation is explicit and is the object of the next Lemma 4.10. In that case we define, for $v,w$ two arbitrary unit orthogonal vectors the quantity
$\beta=\mathbb{E}[\nabla^{2}F(v,w)^{2}],$
which does not depend of the choice of $v,w$ by the invariance by rotation, and the constants
$\beta_{0}=\frac{d}{d+2},\quad\text{and}\quad\gamma=1\pm\sqrt{\frac{2}{d+2}},$
[…]
Lemma 4.9.
Assume that $F$ has an isotropic distribution. For orthogonal unit vectors $v,w\in V$,
$\mathbb{E}[\nabla^{2}F(v,v)^{2}]=3\beta,\quad\mathbb{E}[\nabla^{2}F(v,w)^{2}]=\mathbb{E}[\nabla^{2}F(v,v)\nabla^{2}F(w,w)]=\beta.$
Moreover,
$\beta\geq\beta_{0}$
with equality if and only if $F$ is the isotropic random wave model.
Proof.
In the following, $v,w$ are orthogonal unit vectors in $V$. By symmetry of the derivatives, one has
$\mathbb{E}[\nabla^{2}F(v,w)^{2}]=\partial_{v}\partial_{v}\partial_{w}\partial_{w}q(0)=\mathbb{E}[\nabla^{2}F(v,v)\nabla^{2}F(w,w)].$
Using the invariance by rotation
$E[\nabla^{2}F(v,v)^{2}]$ $=\frac{1}{2}\mathbb{E}\left[\nabla^{2}F\left(\frac{v+w}{\sqrt{2}},\frac{v+w}{\sqrt{2}}\right)^{2}\right]+\frac{1}{2}\mathbb{E}\left[\nabla^{2}F\left(\frac{v-w}{\sqrt{2}},\frac{v-w}{\sqrt{2}}\right)^{2}\right]$
$=\frac{1}{4}\left(\mathbb{E}[\nabla^{2}F(v,v)^{2}]+\mathbb{E}(\nabla^{2}F(w,w)^{2}]\right)+\frac{1}{2}\mathbb{E}[\nabla^{2}F(v,v)\nabla^{2}F(w,w)]+\mathbb{E}[\nabla^{2}F(v,w)^{2}]$
We deduce that
$E[\nabla^{2}F(v,v)^{2}]=\mathbb{E}[\nabla^{2}F(v,v)\nabla^{2}F(w,w)]+2\mathbb{E}[\nabla^{2}F(v,w)^{2}]=3\beta.$
Moreover, a quick computation shows that
$\mathbb{E}\left[\left(F-\frac{1}{d}\Delta F\right)^{2}\right]=\beta\frac{d+2}{d}-1.$
We deduce that $\beta\geq\beta_{0}$ with equality if and only if $F$ is the random wave model. ∎
{mdframed}
[backgroundcolor=white]
Lemma 4.10.
Assume that $F$ has an isotropic distribution. Then the vector
$TF=\frac{1}{\sqrt{2\beta}}\left(\nabla^{2}F-\frac{\gamma}{d}(\Delta F)\mathrm{Id}_{V}\right)$
is a standard Gaussian vector.
Proof.
The constant $\gamma$ must satisfies the identities
@@ 4.2.3 Chaotic expansion under hypothesis (H3)
$0=\mathbb{E}\left[\left(\nabla^{2}F(v,v)-\gamma\frac{\Delta F}{d}\right)\left(\nabla^{2}F(w,w)-\gamma\frac{\Delta f}{d}\right)\right]=\beta\left[1-\frac{2(d+2)}{d}\gamma+\frac{(d+2)}{d}\gamma^{2}\right],$
and
$1=\frac{1}{2\beta}\mathbb{E}\left[\left(\nabla^{2}F(v,v)-\gamma\frac{\Delta F}{d}\right)^{2}\right]=\frac{1}{2}\left[3-\frac{2(d+2)}{d}\gamma+\frac{(d+2)}{d}\gamma^{2}\right].$
We deduce that $\gamma$ is a root of an explicit second degree polynomial, whose roots are given by
$\gamma=1\pm\sqrt{\frac{2}{d+2}}.$
{mdframed}
[backgroundcolor=white]
Lemma 4.11.
Under hypothesis (H3), the spectral measure of $(\nabla F,\widetilde{\nabla^{2}F})$ can be written as
$\mu=\omega\left[2i\pi\xi,S(\xi)\right]\otimes\left[-2i\pi\xi,\overline{S(\xi)}\right],$
where the map $S(\xi)$ is a complex quadratic form in the variable $\xi$. If the distribution of $F$ is isotropic, then the spectral measure of $(\nabla F,TF)$ is
$\mu=\omega\left[2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right]\otimes\left[-2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right].$
Proof.
Both expression follows directly from Lemma 4.4 and Lemma 4.10. ∎
{mdframed}
[backgroundcolor=white]
Lemma 4.12.
Under hypothesis (H3), the second chaotic projection $f_{2}$ has expression
$f_{2}=\alpha\begin{pmatrix}[c|c]-\mathrm{Id}_{V}&0\\ \hline\cr 0&M\end{pmatrix}.$
If the distribution of $F$ is isotropic, then the bilinear map $f_{2}$ has expression, for $v\in V$ and $S$ a symmetric bilinear map
$f_{2}(v,S)=-\alpha\|v\|^{2}+A\operatorname{Tr}(S^{2})+B\operatorname{Tr}(S)^{2}.$
Moreover
$A\left(1-\frac{2\gamma}{2}+\frac{\gamma^{2}}{d}\right)+B\left(1-\gamma\right)^{2}=\frac{2\beta_{0}}{d}\alpha.$
Proof.
The first property is a consequence of the independence between $\nabla F$ and $TF$. As for the second one, we define the constants, for $v,w$ orthogonal unit vectors in $V$
$A=\mathbb{E}\left[f(X)\left(2TF(v,w)^{2}-1\right)\right],\quad B=\mathbb{E}\left[f(X)TF(v,v)TF(w,w)\right]$
and
$C=\mathbb{E}\left[f(X)\left(TF(v,v)^{2}-1\right)\right].$
Using the invariance by rotation one has, following the proof of Lemma 4.10, the identity
$C=A+B.$
Since the distribution of $X$ in invariant by rotation, $f_{2}$ is a quadratic form on the space of symmetric bilinear forms which is invariant by rotation. It must then by a symmetric quadratic form on the eigenvalues, and must then be a linear combination of the quadratic forms $\operatorname{Tr}(S^{2})$ and $\operatorname{Tr}(S)^{2}$. The identification of the coefficient $A$ and $B$ follows by a direct computation of the quadratic form, for instance by fixing an orthogonal basis on $V$.
It also follows from Lemma 4.3 and Equation (4.2.3) that
$d\alpha=\mathbb{E}[f(X)(\|TF\|^{2}-\mathbb{E}[\|TF\|^{2}])]=\frac{d(d-1)}{2}A+dC=\frac{d(d+1)}{2}A+dB.$
The conclusion follows by observing the identities
$A\left(1-\frac{2\gamma}{d}+\frac{\gamma^{2}}{d}\right)+B\left(1-\gamma\right)^{2}=\frac{d+1}{d+2}A+\frac{2}{d+2}B=\frac{2}{d+2}\alpha=\frac{2\beta_{0}}{d}\alpha.$
As in the subsection concerning hypothesis (H1), the fourth chaotic component $f_{4}$ of $f$ can be seen as a symmetric bilinear map on $S^{2}(W^{*})$, which can be decomposed as the direct sum
$S^{2}(W^{*})=S^{2}(V)\oplus S^{2}(S^{2}(V))\oplus\left(V\otimes S^{2}(V)\right).$
{mdframed}
[backgroundcolor=white]
Lemma 4.13.
The subspace $U\otimes S^{2}(V)$ of $S^{2}(W^{*})$ is stable by $f_{4}$. Moreover, the restriction of $f_{4}$ to this subspace, denoted by $f_{4}^{r}$, has expression, for $v\in V$ and $S$ a symmetric bilinear map
$f_{4}^{r}(v\otimes S)=\|v\|^{2}\left(A\operatorname{Tr}(S^{2})+B\operatorname{Tr}(S)^{2}\right),$
where $A,B$ are defined in Lemma 4.12.
Proof.
The fact that the subspace $U\otimes(U\otimes V)$ directly follows from the symmetry of the Jacobian map. The restriction of $f_{4}^{r}$ to this subspace is equal to
$f_{4}^{r}$ $=\mathbb{E}\left[f(X(0))\left(\nabla F(0)\nabla F(0)^{T}-\mathrm{Id}_{V}\right)\otimes\left(TF(0)\otimes TF(0)-\mathrm{Id}_{V}\otimes\mathrm{Id}_{V}\right)\right]$
$=-\mathrm{Id}_{V}\otimes\mathbb{E}\left[f(X(0))\otimes\left(TF(0)\otimes TF(0)-\mathrm{Id}_{V}\otimes\mathrm{Id}_{V}\right)\right]$
$=-\mathrm{Id}_{V}\otimes f_{2}(0,\,\cdot\,).$
The conclusion follows from the expression of $f_{2}$ given by Lemma 4.12. ∎